# Quantifying Emotional Arousal through Pupillary Response: A Novel Approach for Isolating the Luminosity Effect and Predicting Affective States

### v1

@@ Abstract
Researchers have long recognized pupil response as a potential objective indicator of emotional arousal; however, confounding factors, particularly luminosity of stimuli and the ambient environment, have limited its usefulness in detecting emotions. This study presents a new approach to isolate and remove the effect of luminosity on pupil dilation, obtaining the component of pupil dilation due only to emotional arousal.
Our model predicts the pupil size due to luminosity only as a function of the screen luminosity. Our model adapts to individual differences in pupil response to light and different types and configurations of monitors by using a calibration procedure before the experiment. The predicted pupil size has an average correlation with the measured pupil size of 0.76 0.045, an R2 of 0.58 0.680, and a normalized root mean square error (NRMSE) of 0.14 0.014.
Here, we demonstrate that our model can be used simply to calculate emotional arousal. We showed 32 video clips with different content and emotional intensity to 47 participants, who, after each video, reported their level of emotional arousal. We then calculated the pupil size due only to luminosity and subtracted it from the total recorded pupil size, obtaining the component due only to emotional arousal. From the latter, we predicted the arousal of each participant for each video. We obtained an average correlation between predicted and self-reported arousal of 0.65 0.12, an R2 of 0.43 0.12, and an NRMSE of 0.27 0.036. Instead, using the measured pupil size, without subtracting the component due to luminosity, we obtained dramatically worse results: an average correlation between the predicted and self-reported arousal of 0.26 0.15, an R2 of 0.09 0.089, and an NRMSE of 0.42 0.054.
Our results highlight that separating the emotional and luminosity components from pupillary responses is critical to accurately and precisely predicting arousal. Our approach improves the reliability of pupil-based emotional arousal assessment. It opens new avenues for real-time emotion detection in various applications, including human-computer interaction, mental health monitoring, neuromarketing, and affective computing.
Index Terms:
AI (Artificial Intelligence), Pupil Dilation, Pupil Size, Emotion Detection, Arousal, Luminosity Effect, Linear Regression (LR)I Introduction
@@
Emotions are a central component of human life, influencing most, if not all, aspects of cognition and behavior [izard2009emotion]. Arising from complex psychological and physiological processes, they shape how we perceive and interact with the world around us [gross2015emotion], modulating key cognitive functions [tyng2017influences], such as memory, attention, perception, and learning, and shaping our experiences [brosch2013impact, leblanc2015predictable]. Emotions also play an integral role in decision-making [lerner2015emotion, brosch2013impact] and in our social lives through their involvement in empathy, cooperation, and facilitating meaningful interpersonal interactions [van2009emotions]; in workplaces, as they can enhance team dynamics and leadership, contributing to better collaboration and overall productivity [raza2024role, kalraa2024role]; in educational settings [vistorte2024integrating, habibi2025revolutionizing] and in mental health diagnosis and treatment [cooper2020facial].
Accurate emotion detection is a crucial skill across multiple contexts. For example, teachers can observe students’ facial expressions or vocal tones in education to identify emotions like confusion or disengagement [vistorte2024integrating]. This helps them to adjust their teaching methods, ensuring that students remain engaged, especially in remote learning environments where non-verbal cues are more complex to detect [habibi2025revolutionizing]. In mental health, emotion detection plays a key role in diagnosing and treating conditions such as depression, where recognizing patients’ emotional responses is vital for effective care [cooper2020facial].
@@
Emotional changes manifest through various physiological and behavioral indicators, such as heart rate, blood pressure, facial expressions, body language, eye movements, and pupil dilation [shu2018review, galvanophysiological]. Each of these indicators provides valuable insights into emotional states [kassam2013effects], contributing to the development of comprehensive models of human behavior (i.e., Arousal Detection Models) [zeidner2003development]. Accurately detecting emotional arousal has essential applications in fields such as psychology [giannakakis2019review], education [ba2023measuring], human-computer interaction [peter2012emotion], healthcare [tobon2020emotional, dhuheir2021emotion], and artificial intelligence [ballesteros2024facial]. AI-powered emotion detection systems have proven helpful in optimizing emotion detection by analyzing facial expressions [ballesteros2024facial], vocal tones [szelogowski2021emotion], and physiological signals [li2006emotion], enabling more accurate and efficient identification of emotional states. However, AI-powered systems for using pupillary responses to measure emotional arousal have progressed more slowly than those using other physiological and behavioral markers due to the challenge of separating emotional arousal effects from luminosity effects, an issue we address in this study [raiturkar2016decoupling, pan2024effects].
@@
Pupil dilation in response to emotional stimuli is controlled by the autonomic nervous system (ANS) through the interplay of its two branches [lowenstein1950role, partala2003pupil]. The sympathetic nervous system (SNS) triggers pupil dilation (mydriasis) via norepinephrine (a neurotransmitter and hormone), enhancing visual processing during arousal or stress [wu2022ocular]. The parasympathetic nervous system (PNS) regulates pupil constriction (miosis) by acting through the oculomotor nerve and the neurotransmitter acetylcholine. This function is most active during calm or restful states, reducing the amount of light entering the eye and adjusting pupil size for near vision [henderson2020autonomic]. Indeed, previous research shows that pupils dilate within 200 milliseconds of the release of noradrenaline in response to emotional arousal [sirois2014pupillometry, chen2022fatigue, lee2023emotion, alshanskaia2024pupillometry].
It has been well documented that emotional arousal triggers pupil dilation, and ambient luminosity levels also significantly influence pupil size [grujic2024neurobehavioral, pan2024effects, cherng2020background, zhang2019pupil]. Pupils typically constrict in bright conditions and dilate in darker environments [ansari2021evidence], with size varying by up to 50% [spector1990pupils], hindering researchers’ use of pupil dilation as a direct measure of an individual’s emotional state [cherng2020background, mathur2014influences]. Research on emotion detection must account for the impact of luminosity on pupil size, especially in our digital world, where screen brightness and contrast may significantly influence pupil dilation [miranda2018eye]. To reliably use pupil size as an emotional arousal marker, it is crucial to distinguish genuine emotional responses from luminosity-induced changes [binda2014pupil].
Previous research suggests that accounting for luminosity-induced pupil size changes can help to isolate the emotional responses with varying levels of success [guillon2016effects]. Previous research used gray screens to measure baseline pupil diameter [bradley2008pupil, dicriscio2017pupil, kay2022pupillary], but this method has proven insufficient for complex stimuli such as videos with dynamic visual elements [cherng2020background], where changing luminosity and visual elements continually affect pupil responses. In the case of a video, an ideal baseline would be a set of emotionally neutral stimuli of the same luminosity as each video frame, which is challenging to engineer.
In this paper, we propose a more realistic and scalable approach: we directly predicted the component of pupil size influenced by luminosity for each emotional frame instead of including an emotionally neutral frame with the same luminosity repeated for the entire video. This is not the first attempt to use such an approach, with previous studies attempting to disentangle the effects of emotional arousal and light exposure on pupil size for dynamic stimuli such as videos. Our research builds on previous attempts to use such an approach while addressing some of their shortcomings. For example, Nakayama et al. [nakayama2021controlling] developed a hyperbolic model to predict pupil size as a function of the luminosity of the computer screen in a dark laboratory. Other researchers have used simpler models. For example, Tarnowski et al. [tarnowski2020eye] hypothesized a linear relationship between luminosity and pupil size. However, as the relationship between pupil size and luminosity is exponentially decreasing [zhang2019pupil], we developed an exponential model in our previous study [pansara2024towards].
@@
In the present study, we extended and improved our previous work as follows:
@@
First, we tested the model we presented in [pansara2024towards] in a dark laboratory with 10 participants.
@@
Second, we assessed the model’s functionality in a well-lit laboratory environment, testing its robustness across varying luminosity levels.
@@
Third, we calculated the portion of pupil size due to emotional arousal in response to emotional videos, subtracting the portion of pupil size attributable to luminosity from the total pupil size. We asked participants to watch 32 video clips, which varied in emotional content, and report the level of arousal for each of them.
@@
Finally, we tested hyperbolic and linear models, recording the pupil’s size in response to different luminosity levels. We then compared the results with those obtained with our method, demonstrating that our exponential model better describes the relationship between pupil size and luminosity.
Our approach can effectively remove luminosity-induced confounders in pupil-based emotion research, especially in low-light conditions, filling a significant gap. Nakayama and colleagues [nakayama2021controlling] also sought to separate pupil dilation related to emotional arousal from luminosity-induced changes in video content. They also applied their method to emotional images and demonstrated a luminosity and arousal effect on pupil size using ANOVA. We wanted to go one step further: predicting self-reported arousal using pupil size, which was not addressed in Nakayama and colleagues’ work.
Raiturkar et al. [raiturkar2016decoupling] developed a linear model to predict pupil size based on light intensity, allowing for the isolation of emotional arousal by subtracting the effects of light from actual pupil measurements, and applied it to the analysis of emotional videos. However, they did not record self-reported arousal while watching the movie clips.
Where possible, we have applied the existing models mentioned above to our data, with the advantage that we included a comparatively large sample of 47 participants (compared to a maximum of 10 in the studies above), a comparatively rich video collection of 32 emotional video clips, and the fact that we had recorded self-reported arousal for each participant and each video.
In some cases, it was not possible to apply methods from previous studies to our data due to insufficient details regarding the corresponding models. For example, Asano et al. [asano2021neural, asano2021pupil] developed models that considered the temporal evolution of pupil size, which we did not do. Their first model consisted of a linear model and the second of a neural network. We only managed to evaluate the first model but could not assess the neural network model since we could not access their trained neural network. Asano and colleagues [asano2021neural] wrote that the second method performed better than the first one, and they tested both methods with emotional video clips but did not detect self-reported arousal. A more detailed analysis of the differences between our method and those of the other researchers is reported in the Discussion.
In summary, by eliminating confounding luminosity effects, we aimed to improve the accuracy and reliability of pupil-based emotion detection compared to traditional approaches. In the following, we present a robust and easy-to-use algorithm to isolate the effects of brightness on pupil size.
II Methodology
@@
Our approach to developing a model for arousal detection consisted of two study phases:
1.
@@
Development of the Luminosity Effect Prediction Model (LEPM): In this study phase, we developed a model to predict pupil size changes purely caused by luminosity. By subtracting the predicted pupil size from the observed pupil size, we isolated the pupillary response attributable to emotional stimuli.
2.
Development of the Arousal Detection Model (ADM): To show that the LEPM developed in the previous study phase is accurate and has ecological validity, we trained a simple linear model on participants’ pupil size data while watching emotional video clips and used it to predict arousal with and without the LEPM algorithm.
We developed a relatively complex model for the first phase (LEPM) and a very simple one (ADM) for the second phase. Our objective was to demonstrate that the result obtained using the LEPM, i.e., computing the component of pupil size due to arousal, is immediately applicable without knowledge of mathematics or advanced machine learning. As presented here, ADM is deliberately simple, as this study aimed to show the utility of LEPM. Ongoing developments in our lab will build on and extend ADM and its effectiveness in arousal state prediction. In this paper, we wanted to show that our LEPM model allowed us to obtain the component of pupil size due to arousal and that it correlates well with participants’ subjective arousal state.
II-A Development of the Luminosity Effect Prediction Model

### v2

@@ Abstract
Pupil dilation is recognized as an objective indicator of emotional arousal, but confounding factors such as the luminosity of stimuli and the surrounding environment have greatly limited its practical usefulness. This study presents a new approach to isolate and remove the effect of luminosity on pupil dilation. We validated this approach by showing 32 video clips with different content and emotional intensity to 47 participants, who reported their level of emotional arousal after each video. We developed a model capable of predicting the effect of luminosity on pupil size as a function of screen brightness, which adapts to individual physiological differences and different types of monitors through a brief pre-experimental calibration. We thus estimated the pupil size due exclusively to luminosity and subtracted it from the total recorded pupil size, obtaining the component due exclusively to arousal. From the latter, we predicted the arousal of each participant for each video using two models. We first used a simple linear regression model. When we used the luminosity-corrected pupil size, we obtained a correlation between predicted and self-reported arousal of = 0.65 0.12, and of 0.43 0.12. The uncorrected pupil size, instead, showed virtually no predictive power ( = 0.26 0.15, = 0.09 0.089). We then used an Extreme Gradient Boosting model, obtaining even better results in the case of luminosity correction ( = 0.765 0.047, = 0.556 0.085). Our results highlight that separating emotional and luminosity components from pupillary responses is crucial for accurately predicting arousal.
1 Introduction
@@
Emotions are a central component of human life, influencing most, if not all, aspects of cognition and behavior Izard (2009). Arising from complex psychological and physiological processes, they shape how we perceive and interact with the world around us Gross (2015), modulating key cognitive functions Tyng et al. (2017), such as memory, attention, perception, and learning, and shaping our experiences Brosch et al. (2013). Emotions also play an integral role in decision-making Lerner et al. (2015) and in our social lives through their involvement in empathy, cooperation, facilitating meaningful interpersonal interactions Van Kleef (2009), as well as in workplaces, as they can enhance team dynamics and leadership, contributing to better collaboration and overall productivity Kalraa (2024). Recognizing one’s own and others’ emotions is crucial both in educational contexts Vistorte et al. (2024) and in clinical settings, as it is vital for diagnosing and treating mental disorders Cooper et al. (2020).
@@
Emotional changes manifest through various physiological and behavioral indicators, such as heart rate, blood pressure, facial expression, body language, eye movements, and pupil dilation Shu et al. (2018). Each of these indicators provides valuable insights into emotional states Kassam and Mendes (2013), supporting the development of comprehensive models of emotional and behavioral processes Zeidner et al. (2003). Accurately detecting emotional arousal has valuable applications in fields such as psychology Giannakakis et al. (2019), education Ba and Hu (2023), human-computer interaction Peter and Urban (2012), healthcare Dhuheir et al. (2021), and artificial intelligence Ballesteros et al. (2024). AI-powered emotion detection systems have been proven helpful for optimizing emotion detection by analyzing facial expressions Ballesteros et al. (2024), vocal tone Szelogowski (2021), and physiological signals Wang and Wang (2025), enabling more accurate and efficient identification of emotional states.
@@
AI-powered systems using pupillary responses to measure emotional arousal have progressed more slowly than those using other physiological and behavioral markers due to the challenge of separating emotional arousal effects from luminosity effects, an issue we address in these studies Raiturkar et al. (2016); Pan et al. (2024). Foundational work by Bradley et al. established that while an initial light reflex occurs, the pupil’s subsequent response is larger when viewing emotionally arousing stimuli, regardless of valence, a change mediated by sympathetic activity Bradley et al. (2008). Pupil dilation in response to emotional stimuli is indeed controlled by the autonomic nervous system (ANS) through the interplay of its two branches Lowenstein and Loewenfeld (1950); Partala and Surakka (2003). The sympathetic nervous system (SNS) triggers pupil dilation (mydriasis) via the neurotransmitter norepinephrine, enhancing visual processing during arousal or stress Wu et al. (2022). The parasympathetic nervous system (PNS) regulates pupil constriction (miosis) by acting through the oculomotor nerve and the neurotransmitter acetylcholine. This function is most active during calm or restful states, reducing the amount of light entering the eye and adjusting pupil size for near vision Henderson and Spies (2020). Previous research shows that pupils often dilate within 200 milliseconds of the release of noradrenaline in response to emotional arousal Sirois and Brisson (2014).
In addition to emotional arousal, ambient light levels also significantly influence pupil size. Pan et al. (2024). Pupils typically constrict in bright conditions and dilate in darker environments Mathur et al. (2014), with size varying by up to 50% Spector (1990), hindering researchers’ use of pupil dilation as a direct measure of an individual’s emotional state Cherng et al. (2020). A central challenge in pupillometry is, therefore, delineating the observed pupil size signal into its dominant Pupillary Light Response (PLR) term () and its psychologically relevant Psychosensory Pupil Response (PPR), or arousal term (). As Mathôt et al. and Strauch et al. detail, this challenge stems from the pupil acting as a “final common pathway” (or integrated readout) for both the light reflex and the emotional response Mathôt (2018); Strauch et al. (2022). Therefore, explicit modeling and removal of the luminance term is a necessary step for reliable arousal inference. Establishing robust pre-processing and controlling for these luminance confounds are the “gold standards” for valid inference Kret and Sjak-Shie (2019); Mathôt and Vilotijević (2022). However, as the field moves toward more advanced dynamic modeling of pupil data Fink et al. (2023), it has become increasingly clear that traditional techniques like subtractive baseline correction are often insufficient for dynamic visual stimuli where luminance varies continuously.
Foundational work established the first mathematical models for predicting pupil size under controlled light. The unified formula of Watson and Yellott provides a parametric characterization of the steady-state PLR based on luminance, field size, and age Watson and Yellott (2012). On the high-precision end of the engineering spectrum, Zhang et al. proposed a model based on Spatially Weighted Corneal Flux Density, a high-precision metric used to estimate pupil size by accounting for the non-uniform sensitivity of the retina to light Zhang et al. (2019). Extending this, Huang et al. showed that PLR is sensitive to the spatial and chromatic distributions of light across the retina, indicating that accurate modeling requires spatially
aware methods Huang et al. (2023).
Nakayama et al. explored several mathematical formulations to express the relationship between pupil size and brightness, including power-log and exponential functions, before ultimately selecting a hyperbolic model to predict pupil size as a function of screen luminosity in a dark laboratory environment Nakayama et al. (2021). This model is personalized such that the specific coefficients of the hyperbolic function are determined experimentally for each individual to account for personal physiological differences in light adaptation. Although they used subjective evaluation scores and task conditions to validate the model’s success in isolating signals related to mental activity, they did not utilize the resulting residual to predict participants’ self-reported arousal.
An alternative research tradition treats the pupil as a dynamic system whose neural inputs are convolved with an impulse-response function. Wierda et al. used deconvolution to uncover fast attentional dynamics from pupil traces Wierda et al. (2012). This was refined by McCloy et al. who improved the alignment between model predictions and behavioral events by estimating the pupil impulse response McCloy et al. (2016). Korn and Bach formalized this using the General Linear Model (GLM) to jointly estimate inputs from light and cognitive events Korn and Bach (2016). However, these studies use a temporal attention task with sequential, isoluminant, and discrete stimuli (letters) or, more generally, controlled cognitive events, i.e. stimuli in which the onset of the cognitive event is precisely known. In many laboratory conditions where naturalistic stimuli are used (e.g., emotionally charged videos), the arousal events and their timing are ambiguous, which violates the requirement for discrete and convoluted input signals. This framework was extended by Napieralski and Rynkiewicz to explicitly model the PLR for the purpose of decoupling light-driven dilation from other factors Napieralski and Rynkiewicz (2019). However, that study did not include validation where both factors-dynamic luminosity and emotional arousal are present simultaneously. In summary, the dynamic models presented in the above-mentioned studies are powerful for controlled, event-based experiments, but do not extrapolate to continuous naturalistic content like videos where luminance varies unpredictably.
Asano et al. developed dynamic reaction models using layered neural networks to simulate pupil behavior during continuous temporal changes in brightness, which proved more effective than simple linear models for dynamic video viewing Asano et al. (2021a). Although this approach successfully reconstructed the light reflex to isolate signals related to mental activity induced by audio instructions, the authors did not utilize the model to predict participants’ self-reported arousal. Recently, Fanourakis and Chanel improved a highly detailed biologically plausible dynamic model of the PLR, devised by Pamplona et al., using a numerical solver to attenuate the dynamic light response during continuous screen viewing Fanourakis and Chanel (2022); Pamplona et al. (2009). However, they used this complex model in a generalized, group-level fitting procedure, i.e., the model’s parameters (e.g., sphincter/dilator tension, time constants) were not determined by a separate, dedicated calibration task for each individual subject Fanourakis and Chanel (2022). Recently, Cai et al. introduced Open-DPSM, a convolutional toolkit that models responses to luminance, contrast, and spatial weighting Cai et al. (2024). While Open-DPSM represents the state-of-the-art in sensory modeling, it requires specialized hardware (photometers) for physical luminance calibration and complex gaze-contingent setups. Furthermore, as a sensory-driven model, it has not been validated on emotional or affective stimuli.
For applications requiring continuous processing, data-driven signal-processing methods have been employed. Gao et al. used an Adaptive Interference Canceler (AIC) driven by a photodiode-measured ambient light signal to isolate an affective pupil response Gao et al. (2009). Soleymani et al. provided a multimodal database and conducted emotion classification using standard statistical and spectral features (e.g., average, standard deviation, and Hippus effect) derived from the pupil signal, achieving a 63.5 % classification rate for arousal (low, medium, high) using eye gaze features Soleymani et al. (2012). This approach relied on standard filtering and feature selection but lacked an explicit computational model for dynamic PLR removal, reinforcing the need for targeted PLR mitigation. Further efforts to isolate affective signals in videos led to the “decoupling” framework proposed by Raiturkar et al. Raiturkar et al. (2016). They utilized a linear model of the PLR to predict pupil size based on frame luminance, treating the residual between measured and predicted size as an indicator of emotional arousal. While this laid the groundwork for signal separation, the model’s performance was primarily validated through emotional event peaks identified by independent coders rather than continuous participant self-reports. Later in this article, we will explain why participants’ self-assessments allow us to build models with greater predictive power than the evaluation of stimuli by independent encoders. More recently, Tarnowski et al. utilized a subject-specific linear regression to remove the effect of movie luminance from the pupil signal before performing emotion classification Tarnowski et al. (2020). Although this approach considered individualized modeling and achieved a high accuracy of 80% on a 3-class classification task, it remains interesting to consider how a linear suppression model would perform when predicting continuous arousal. A possible limitation of this approach is that the model relies on a linear approximation for the inherently non-linear, exponential-like pupillary light reflex.
A recent critical finding is that the interaction between arousal and light is multiplicative, not additive Pan et al. (2022). Pan et al. first demonstrated this for emotional arousal Pan et al. (2022). Crucially, Pan et al. later showed that emotional arousal interacts differently, with the strongest effects occurring at very low luminances (), confirming that arousal effects are not monolithic Pan et al. (2024). These studies establish the non-linear structure required for accurate modeling before PPR can be isolated. Also, Cherng et al. demonstrated that background luminance levels significantly modulate affective responses, with greater dilation typically observed under darker conditions due to competing activations of the sympathetic and parasympathetic nervous systems Cherng et al. (2020). Conversely, the models that assume a linear relationship between arousal and luminosity presented in this introduction have a good predictive power, probably because under normal experimental conditions the variation in luminosity is small enough to allow for a linear approximation with limited error. However, it is clear that this interaction highlights the need for a calibration process capable of accurately mapping the non-linear relationship between light and pupil size for each individual and the use of non-linear models.
In conclusion, previous research confirms that PLR modeling is essential for arousal inference from pupil dilation, that the light-arousal interaction is non-linear, and that PLR suppression improves emotion inference. However, no prior computational method provides a single solution that simultaneously offers per-subject calibration, continuous frame-wise processing compatible with video, no need for discrete stimulus timing, and with direct validation of the arousal residual against self-report. The present study introduces such an approach by fitting an individual luminance-pupil function and subtracting its prediction to recover an arousal-sensitive residual that explains self-reported arousal substantially better than raw pupil size. We recently established the foundation for this methodology in Pansara et al., where we demonstrated the technical feasibility of using a personalized non-linear model to isolate the component during video viewing Pansara et al. (2024). However, that preliminary work was limited to the predictive accuracy of the light-effect model itself and did not include removing it to isolate emotional arousal effects. Furthermore, that study relied on a relatively small participant sample. The current work extends that effort by applying the model to a larger and more diverse sample, including participants recruited outside the university population, and shifting from signal isolation to the direct prediction of continuous affective states, providing the first real-time validation against participants’ self-reports of arousal.
@@
In particular, in the present study, we extended and improved our previous work as follows:
@@
First, we tested the model we presented in Pansara et al. (2024) in a dark laboratory with 10 participants.
@@
Second, we assessed the model’s functionality and tested its robustness across varying luminosity levels in laboratory conditions.
@@
Third, we calculated the portion of pupil size change due to emotional arousal of 47 participants watching 32 emotional video clips, by subtracting the portion of pupil size attributable to luminosity from the total pupil size.
@@
Finally, we implemented a machine learning framework using Gradient Boosting to validate the physiological relevance of our isolated residual. After removing the luminosity-driven component, we used the resulting arousal-sensitive residual to predict participants’ self-reported emotional arousal.
The primary goal of this study was to provide a realistic and scalable tool for researchers – particularly in the field of psychology – conducting laboratory experiments with naturalistic video stimuli, and without the need to use specialized hardware such as a photometer. Our goal was to address the ambient light conditions and dynamic screen luminance found in typical experimental settings, rather than addressing the general problem of pupil size variation with luminosity (e.g., with extreme luminosity variations in open environments). These are fundamental steps towards providing a tool to assess emotional arousal from pupil dilation in dynamic screen-based conditions, where luminosity levels can change almost constantly.
2 Methodology
@@
Our approach to developing a model for arousal detection consisted of three study phases:
A.
@@
Development of the Luminosity Effect Prediction Model (LEPM): In this study phase, we developed a model to predict pupil size changes purely caused by luminosity and remove its effect.
B.
Development of the Arousal Detection Model (ADM): We trained a linear model to predict emotional arousal from participants’ pupil size data while they watched emotional video clips. To do this, we considered two conditions: in the first, we corrected the measured pupil size by removing the effect of luminosity using the LEPM model; in the second, we left the signal uncorrected. We then trained a linear regression model in both conditions using the participants’ self-reported arousal as the target variable. Finally, we compared the results in the two conditions. At this stage, we deliberately chose a simple linear regression model to maximize accessibility and ease of use for researchers, while maintaining high predictive performance.
C.
Development of the Gradient Boost Regression-based Self-reported Arousal Prediction Model (GBR-SAPM): To achieve even better performance (than the ADM) in predicting arousal from pupil dilation, we utilized an Extreme Gradient Boosting (XGBoost) regressor to predict continuous arousal states, employing statistical features extracted from both raw and luminosity-corrected pupil signals. To ensure robust, subject-independent generalization, the model was evaluated using a nested Leave-One-Participant-Out (LOPO) cross-validation framework. This approach allows the model to capture complex non-linear physiological dynamics while minimizing the risk of overfitting to individual participants.
2.1 Development of the Luminosity Effect Prediction Model